OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 2x - 3y = -18 PLEASE PLEASE PLEASE HELP !

OpenStudy (ivettef365):

just substitute the x for 0 to find y and then the y for 0 to find x

OpenStudy (ivettef365):

what is x when y=0 you have 2x - 3(0) = -18 (solve for x)

OpenStudy (anonymous):

To find the x-intercepts, let y = 0 2x - 3y = -18 2x - 3(0) = -18 2x - 0 = -18 . . . ----------------------------------------------------------------------- To find the y-intercepts, let x = 0 2x - 3y = -18 2(0) - 3y = -18 0 - 3y = -18 . . .

OpenStudy (anonymous):

Answer is?

OpenStudy (anonymous):

To find the x-intercepts, let y = 0 2x - 3y = -18 2x - 3(0) = -18 2x - 0 = -18 2x = -18 --- ---- 2 2 x = ?

OpenStudy (anonymous):

To find the y-intercepts, let x = 0 2x - 3y = -18 2(0) - 3y = -18 0 - 3y = -18 -3y = -18 ---- ---- -3 -3 y = ?